# Generalized Multi-Objective Reinforcement Learning with Envelope Updates in URLLC-enabled Vehicular Networks

### v1

@@ III-B3 Rewards
where $\text{WR}_{i^{*},j,t}$ is the achievable data rate compute by (6) and $\xi^{j}_{t}$ is the HO probability of AV $j$ computed by dividing the number of HOs accounted until the current time $t$ by the time duration of previous time slots in the episode [^3] Note that $c_{1}\dots c_{5}$ are the weights to set the priority of each term. For instance, $c_{2}$ needs to be sufficiently large compared to other coefficients for collision avoidance. The highest penalty applies to vehicle collision..
@@ IV Proposed Single-Policy and Multi-Policy MORL Algorithms
IV-A Single-Policy MORL Algorithms
Given a set of preferences in MORL problems, single policy algorithms aim to scalarize the reward value to determine the best policy, considering the relative priorities assigned to competing objectives. We explore two DRL methods: DQN [1] and DDQN [25] for MORL, each of which employs a neural network parameterized by $\boldsymbol{\theta}_{t}$ (in each time step $t$) to approximate the $Q$-value function for a state-action pair, i.e., $Q(s_{t},a_{t};\boldsymbol{\theta}_{t})$. After taking action $a_{t}$ in state $s_{t}$ and receiving instant reward $r_{t+1}$, we can formulate a target $Q$ function ${\hat{Q}}(s_{t},a_{t})$ as in Eq. 16, which is used to optimize the neural network $\boldsymbol{\theta}_{t}$ using gradient descent, as given in [25],
@@ IV-B Multi-Policy Envelope MORL Algorithm
Our proposed MO-DDQN-envelope algorithm is designed to learn a spectrum of Pareto optimal policies simultaneously (as defined in Section III-A2) in a preference space ${\Omega}$, as illustrated in Fig. 2. Different from the Envelope-MOQ model in [24], the proposed MO-DDQN envelope algorithm incorporates DDQN instead of using the original REINFORCE algorithm to improve convergence and sample training efficiency.
@@ IV-B1 Bellman Operation with Optimal Filter
In the context described in Section III-A and referenced by [24], the expected discounted return under a policy $\pi$ is defined as $\mathbf{Q}_{\pi}(s,a,\boldsymbol{\omega})=\mathbb{E}_{\pi}\left[\mathbf{r}(s_{t},a_{t})+\gamma\mathbf{Q}_{\pi}(s_{t-1},a_{t-1},\boldsymbol{\omega})\right]$. Yang et al. in [24] further introduces the concept of an optimal filter $\mathcal{H}$ [^4] The optimal filter $\mathcal{H}$ is instrumental in solving the convex envelope of PPF, which represents the current solution frontier. This process is key in optimizing the Q-function, $\mathbf{Q}_{\pi}$ for a given state $s$ and preference weights $\boldsymbol{\omega}$., which is applied to $\mathbf{Q}_{\pi}(s,a,\boldsymbol{\omega})$ to obtain $(\mathcal{H}\mathbf{Q})_{\pi}(s,a,\boldsymbol{\omega})=\arg_{Q}\sup_{a\in\mathcal{A},\boldsymbol{\omega}^{\prime}\in\Omega}\mathbf{Q}_{\pi}(s,a,\boldsymbol{\omega}^{\prime})$. Here, $\boldsymbol{\omega}$ is optimized through a process that balances preference between objectives, i.e., transportation and telecommunication. The $\arg_{Q}$ represents a multi-objective supremum value, ensuring that $(a,\boldsymbol{\omega}^{\prime})$ achieves the maximum supremum across actions in space $\mathcal{A}$ and states $\boldsymbol{\omega}^{\prime}$ within the space $\Omega$. Consequently, we can streamline (17) to focus the optimization on actions solely dependent on $\mathcal{H}$. The MO optimality operator can thus be defined as:
@@
Based on the specific Bellman operator and the optimal filter $\mathcal{H}$, we maintain the convex envelope $\sup_{\boldsymbol{\omega}^{\prime}}\boldsymbol{\omega}^{T}\mathbf{Q}_{\pi}(s,a,\boldsymbol{\omega}^{\prime})$, corresponding to new preference weights for optimal MO rewards. These rewards may not be optimized by other past preference weights during training. Unlike benchmark approaches discussed in Section IV-A, single policy scalarized updates fail to optimize the scalar utility for acquiring the optimal solution with varying $\boldsymbol{\omega}$, due to their inability to leverage information from $\max_{a}\mathbf{Q}_{\pi}(s,a,\boldsymbol{\omega})$.
@@ IV-B3 Homotopy Optimization
for $\forall z\in[1,N_{\mathcal{T}}]$ and $\forall g\in[1,N_{\omega}]$. Finding the optimal preference weight $\boldsymbol{\omega}^{\prime}$ in $\Omega$ can be an NP-hard problem due to the size and complexity of $\Omega$. Instead, finding the optimal preference $\boldsymbol{\omega}^{\prime}$ in $\mathcal{W}$ is feasible. By replay sampling transition $(s_{t},a_{t},\mathbf{r}_{t},s_{t+1})$ across $N_{\mathcal{T}}$ transitions, we acquire the empirical estimate target function over new state $s_{t+1}$ as:
@@
In homotopy optimization, $\lambda_{t}$ gradually increases from 0 to 1 through balance weight path $p_{\lambda}$, which adjusts the equilibrium between $\mathcal{L}^{A}(\boldsymbol{\theta}_{t})$ and $\mathcal{L}^{B}(\boldsymbol{\theta}_{t})$. Loss function $\mathcal{L}(\boldsymbol{\theta}_{t})$ smoothly shift from $\mathcal{L}^{A}(\boldsymbol{\theta}_{t})$ to $\mathcal{L}^{B}(\boldsymbol{\theta}_{t})$ to ensure achieving first accuracy $\mathbf{Q}$ optimization, then smoothing utility provided by auxiliary. We first trying to reduce the discrepancy between target and estimate $\mathbf{Q}$ value as $(\hat{\mathbf{Q}}(s_{t},a_{t},\omega_{t};\boldsymbol{\theta}_{t})-\mathbf{Q}(s_{t},a_{t},\omega_{t};\boldsymbol{\theta}_{t})$ and then taking gradient descent $\nabla_{\boldsymbol{\theta}_{t}}$ for estimate $\mathbf{Q}$ value to adjust direction to reduce MSE. Consequently, parameter for MO-DDQN-envelope will be updated as,
@@ IV-C Complexity Analysis
Therefore, the overall time complexity for MO-DDQN-envelope can be expressed as $\mathcal{O}(M_{1}\cdot T_{\text{hl}}\cdot|\mathcal{A}|+N_{\mathcal{T}}\cdot N_{\omega}+\sum_{e=1}^{E+1}(N_{e-1}\cdot N_{e})+|\mathcal{S}|\cdot|\mathcal{A}|)$.
@@ V-A Simulation Environment
We utilize a Multi-Layer Perceptron (MLP) to construct the $Q_{\theta}$ neural network in the training and evaluation phases. Our architecture comprises of three-layer FNNs, with each layer hosting 128 neurons. Following these 3 layers of FNNs, we apply the ReLU activation function. The output layer of the target policy network $Q_{\theta}^{\prime}$ employs a sigmoid function to constrain the range of actions. Apart from the input and output layers, the MO evaluation network $Q_{\theta}$ and the target policy network $Q_{\theta}^{\prime}$ share identical architectures and utilize the same activation functions. Simulation parameters are detailed in Table I unless otherwise specified.
Table I: Values of system parameters in experiments
@@ V-B Baselines and Evaluation Metrics
To comprehensively evaluate the performance of MO-DDQN-envelope, we implement four baseline algorithms for comparison: MO-DQN, MO-DDQN (as discussed in Section IV-A), and two additional MORL algorithms, namely MO-dueling-DDQN and MO-PPO. MO-dueling-DDQN and MO-PPO are variations of MO-DDQN, employing single-policy multi-objective dueling DDQN and Proximal Policy Optimization (PPO) algorithms instead of DDQN, respectively. These algorithms are developed specifically for performance evaluation. It is noteworthy that dueling-DDQN and PPO are extensions of single-policy single-objective RL algorithms proposed in [33] and [34], respectively.
To evaluate the performance, we consider the following metrics. Assume episode $e$ ends on time step $T_{e}$. we define (1) total transportation reward: $R^{tran}_{e}=\mathbb{E}_{j\in M_{1}}[\sum_{t=1}^{T_{e}}r^{\mathrm{j,tran}}_{t}]$. (2) total communication reward: $R^{tele}_{e}=\mathbb{E}_{j\in M_{1}}[\sum_{t=1}^{T_{e}}r^{\mathrm{j,tele}}_{t}]$. (3) collision rate: $\delta_{e}=1-\frac{T_{e}}{T_{hl}}$ (4) HOs Probability: $\xi_{e}=\mathbb{E}_{j\in M_{1}}[\xi^{j}_{T_{e}}]$. where $r^{\mathrm{j,tran}}_{t}$ and $r^{\mathrm{j,tele}}_{t}$ are instantaneous transportation and telecommunication rewards specified in equations (20) and (21), respectively. $T_{hl}$ represents the horizon limit of each episode. $\xi^{j}_{T_{e}}$ denotes HOs probability by the end of each episode ($T_{e}$’s step) defined in equation (21).

### v2

@@ III-A2 MOMDPs
From Fig. 2(a), the Pareto frontier includes points A to H, encapsulating some local concave regions. Grey points indicate non-optimal solutions that do not belong to the CCS in terms of transportation and telecommunication objective. In Fig. 2(b), linear preferences are used to select the optimal solution from the CCS based on the highest utility. This utility corresponds to the projection length of each point onto the preference vector $\omega_{i}\in\Omega$. Blue arrows represent different linear preferences reflecting trade-offs between transportation and telecommunication objectives. Among the possible returns, point C achieves a higher cumulative utility compared to points D, I, and H when projected along the preference vector represented by the solid line $\omega_{i}$. $\omega_{i}\cdot\hat{r}_{C}>\omega_{i}\cdot\hat{r}_{D}>\omega_{i}\cdot\hat{r}_{I}>\omega_{i}\cdot\hat{r}_{H}$. the optimal solution is the point in the CCS with the largest projection along the preference vector $\omega_{i}$
@@ III-B3 Rewards
where $c_{5}$ is the coefficient to scalarlize weighted datarate. $\text{WR}_{i^{*},j,t}$ is the achievable data rate compute by (6) and $\xi^{j}_{t}$ is the HO probability of AV $j$ computed by dividing the number of HOs accounted until the current time $t$ by the time duration of previous time slots in the episode [^3] Note that $c_{1}\dots c_{5}$ are the weights to set the priority of each term. For instance, $c_{2}$ needs to be sufficiently large compared to other coefficients for collision avoidance. The highest penalty applies to vehicle collision..
@@ IV Proposed Single-Policy and Multi-Policy MORL Algorithms
IV-A MORL Solution with Predefined Preferences
Given a set of preferences in MORL problems, single policy algorithms aim to scalarize the reward value to determine the best policy, considering the relative priorities assigned to competing objectives. We explore two DRL methods: DQN [1] and DDQN [32] for MORL, each of which employs a neural network parameterized by $\boldsymbol{\theta}_{t}$ (in each time step $t$) to approximate the $Q$-value function for a state-action pair, i.e., $Q(s_{t},a_{t};\boldsymbol{\theta}_{t})$. After taking action $a_{t}$ in state $s_{t}$ and receiving instant reward $r_{t+1}$, we can formulate a target $Q$ function ${\hat{Q}}(s_{t},a_{t})$ as in Eq. (13), which is used to optimize the neural network $\boldsymbol{\theta}_{t}$ using gradient descent, as given in [32],
@@ IV-B Multi-Policy Envelope MORL Algorithm
Our proposed MO-DDQN-Envelope algorithm is designed to learn a spectrum of Pareto-optimal policies simultaneously within a preference space ${\Omega}$, as described in Section IV-A and illustrated in Fig. 3. Unlike the Envelope-MOQ model proposed in [31], which employs the REINFORCE algorithm, our MO-DDQN-Envelope algorithm incorporates DDQN to enhance both convergence stability and sample training efficiency. REINFORCE, as a policy gradient method, often suffers from high variance in complex environments like RF-Thz-Highway-Env, leading to unstable updates. In contrast, DDQN leverages temporal difference (TD) learning, which reduces variance and mitigates the overestimation bias. Furthermore, DDQN employs replay experience to enable the agent to learn from past experiences multiple times, improving sample efficiency, and uses target networks to stabilize learning by ensuring consistent targets during updates.
@@ IV-B1 Bellman Operation with Optimal Filter
In the context described in Section III-A and referenced by [31], the expected discounted return under a policy $\pi$ is defined as $\mathbf{Q}_{\pi}(s,a,\boldsymbol{\omega})=\mathbb{E}_{\pi}\left[\mathbf{r}(s_{t},a_{t})+\gamma\mathbf{Q}_{\pi}(s_{t-1},a_{t-1},\boldsymbol{\omega})\right]$. Yang et al. in [31] further introduces the concept of an optimal filter $\mathcal{H}$ [^4] The optimal filter $\mathcal{H}$ is instrumental in solving the convex envelope of PPF, which represents the current solution frontier. This process is key in optimizing the Q-function, $\mathbf{Q}_{\pi}$ for a given state $s$ and preference weights $\boldsymbol{\omega}$., which is applied to $\mathbf{Q}_{\pi}(s,a,\boldsymbol{\omega})$ to obtain $(\mathcal{H}\mathbf{Q})_{\pi}(s,a,\boldsymbol{\omega})=\arg_{Q}\sup_{a\in\mathcal{A},\boldsymbol{\omega}^{\prime}\in\Omega}\mathbf{Q}_{\pi}(s,a,\boldsymbol{\omega}^{\prime})$. The $\arg_{Q}$ represents a multi-objective supremum value, ensuring that $(a,\boldsymbol{\omega}^{\prime})$ achieves the maximum supremum across actions in space $\mathcal{A}$ and states $\boldsymbol{\omega}^{\prime}$ within the space $\Omega$. Consequently, we utilize (14) to focus the optimization on actions solely dependent on $\mathcal{H}$. The MO optimality operator can thus be defined as:
@@
During the training phase, $\boldsymbol{\omega}$ values are collected through free exploration of the environment. As can be seen in Fig. 2, $\omega_{i}$ specifies preference weights to balance objectives. In the initial free exploration phase, $\omega_{i}$ uniformly samples from the sampling space $\Omega$. Uniform sampling ensures that the exploration covers a broad range of trade-offs, increasing the likelihood of finding Pareto-optimal solutions across the entire objective space.
Unlike scalarized single-policy approaches discussed in Section IV-A, which fail to adapt the scalar utility across varying $\boldsymbol{\omega}$, the convex envelope formulation explicitly leverages the supremum operator to optimize across all possible actions $a$ and preference vectors $\boldsymbol{\omega}^{\prime}$.
@@ IV-B3 Homotopy Optimization
for $\forall z\in[1,N_{\mathcal{T}}]$ and $\forall g\in[1,N_{\omega}]$. Finding the optimal preference weight $\boldsymbol{\omega}^{\prime}$ in $\Omega$ can be an NP-hard problem due to the size and complexity of $\Omega$. Instead, finding the optimal preference $\boldsymbol{\omega}^{\prime}$ in $\mathcal{W}$ is feasible. In contrast to $\Omega$, $\mathcal{W}$ contains only those preferences that align with optimal solutions. This is because $\mathcal{W}$ corresponds to the convex boundary of the Pareto frontier, where the utility projection $\boldsymbol{\omega}^{\top}\mathbf{Q}(s,a,\boldsymbol{\omega}^{\prime})$ achieves its maximum for some preference $\boldsymbol{\omega}^{\prime}\in\Omega$. Consequently, $\mathcal{W}$ serves as a reduced yet comprehensive subset of $\Omega$, facilitating efficient exploration and optimization without excluding any optimal solutions. From a computational perspective, the use of $\mathcal{W}$ simplifies the optimization process. By maintaining the envelope $\sup_{\boldsymbol{\omega}^{\prime}}\boldsymbol{\omega}^{\top}\mathbf{Q}(s,a,\boldsymbol{\omega}^{\prime})$, the algorithm effectively focuses on preferences that yield maximal utility, avoiding unnecessary evaluations in suboptimal regions of $\Omega$. This ensures that the updated policies and preferences are always aligned with the optimal frontier. Furthermore, the iterative refinement of $\mathcal{W}$ through envelope updates enables the model to integrate information from previously explored trajectories stored in the preference pool $D_{\omega}$. This process results in faster convergence and improved sample efficiency compared to directly exploring the entire space $\Omega$. By focusing optimization within $\mathcal{W}$, the approach avoids unnecessary evaluations in suboptimal regions of $\Omega$, thereby improving computational efficiency.
By replay sampling transition $(s_{t},a_{t},\mathbf{r}_{t},s_{t+1})$ across $N_{\mathcal{T}}$ transitions, we acquire the empirical estimate target function over new state $s_{t+1}$ as:
@@
We progressively increase the value of $\lambda_{t}$ from 00 to $1$ throughout the training process, thereby transitioning the loss function from $\mathcal{L}^{A}(\boldsymbol{\theta}_{t})$ to $\mathcal{L}^{B}(\boldsymbol{\theta}_{t})$. During training, $\lambda_{t}$ continuously evolves within the range $10<\lambda_{t}<$ 1, ensuring a gradual shift that balances the contributions of the two objectives as the training advances [37, 31]. This method called homotopy optimization [37] is effective since for each update step, it uses the optimization result from the previous step as the initial guess. In the envelope deep MORL algorithm, $\mathcal{L}^{A}(\boldsymbol{\theta}_{t})$ first ensures the prediction of $\mathrm{Q}$ is close to any real expected total reward, though not necessarily optimal. And, then $\mathcal{L}^{B}(\boldsymbol{\theta}_{t})$ can pull the current guess along the direction with better utility. As depicted in Figure 4, the MSE loss $\mathcal{L}^{A}(\boldsymbol{\theta}_{t})$ is difficult to optimize since there are many local minima over its landscape. Although the loss of the value metric $\mathcal{L}^{B}(\boldsymbol{\theta}_{t})$ has fewer local minima, it is also difficult for optimization since there are many vectors $\mathrm{Q}$ minimizing value metric $\omega_{t}^{T}\cdot|\hat{\mathbf{Q}}(s_{t},a_{t},\omega_{t};\boldsymbol{\theta}_{t}^{\prime})-\mathbf{Q}(s_{t},a_{t},\omega_{t};\boldsymbol{\theta}_{t})|$, making $\mathcal{L}^{B}(\boldsymbol{\theta}_{t})$ is flat. The homotopy path connecting $\mathcal{L}^{A}(\boldsymbol{\theta}_{t})$ and $\mathcal{L}^{B}(\boldsymbol{\theta}_{t})$ provide better opportunities to find the optimal global parameters $\mathcal{L}(\boldsymbol{\theta}_{t})$
Figure 4: An explanation for homotopy optimization method used in the MO-DDQN-Envelope
We first trying to reduce the discrepancy between target and estimate $\mathbf{Q}$ value as $(\hat{\mathbf{Q}}(s_{t},a_{t},\omega_{t};\boldsymbol{\theta}_{t})-\mathbf{Q}(s_{t},a_{t},\omega_{t};\boldsymbol{\theta}_{t})$ and then taking gradient descent $\nabla_{\boldsymbol{\theta}_{t}}$ for estimate $\mathbf{Q}$ value to adjust direction to reduce MSE. Consequently, parameter for MO-DDQN-envelope will be updated as,
@@ IV-C1 Training Complexity Analysis
Therefore, the overall training complexity for MO-DDQN-envelope on 1 episode can be expressed as $\mathcal{O}(M_{1}\cdot T_{\text{hl}}\cdot|\mathcal{A}|+N_{\mathcal{T}}\cdot N_{\omega}+\sum_{e=1}^{E+1}(N_{e-1}\cdot N_{e})+|\mathcal{S}|\cdot|\mathcal{A}|)$.
IV-C2 Evaluation Complexity Analysis
Assume we have $N_{l}$ layers in the neural network, each containing $N_{u}$ neurons. According to Section III-B, the neural network processes $|\mathcal{S}|$ inputs from the state space and produces $|\mathcal{A}|$ possible actions as output. The overall complexity for a forward pass through the network is given by:
$\mathcal{O}(|\mathcal{S}|+|\mathcal{A}|+N_{u}^{N_{l}})\approx\mathcal{O}(N_{u}^{N_{l}}),$
and $N_{u}^{N_{l}}$, the computation for the hidden layers, dominates the complexity. For evaluation, we assume a single step is implemented with a constant-time complexity of $\mathcal{O}(1)$. If there are $M_{1}$ target AVs and each travels for at most $T_{hl}$ time steps in one episode, the total number of steps in an episode is $M_{1}\cdot T_{hl}$. Thus, the overall complexity for one simulation episode is:
$\mathcal{O}(M_{1}\cdot T_{hl}\cdot(N_{u}^{N_{l}}+\mathcal{O}(1)))\approx\mathcal{O}(M_{1}\cdot T_{hl}\cdot N_{u}^{N_{l}}),$
where the constant $\mathcal{O}(1)$ is negligible compared to the forward pass complexity.
@@ V-A Simulation Environment
We utilize a Multi-Layer Perceptron (MLP) to construct the $Q_{\theta}$ neural network for the training and evaluation phases. The architecture consists of three fully connected layers (FNNs), each containing 128 neurons. Following these layers, we apply the ReLU activation function to improve training efficiency. The output layer of the target policy network $Q_{\theta}^{\prime}$ employs a sigmoid activation function to constrain the output range of actions. This ensures that the action values remain within a feasible and interpretable range, which is critical for MOO decision-making tasks. This architecture was chosen after a comprehensive hyperparameter tuning process. Simulation parameters are detailed in Table II unless otherwise specified.
@@
Table II: Values of system parameters in experiments
We evaluated all methods across five test instances, varying key variables: desired minimum and maximum longitudinal velocities ($v_{\text{min}},v_{\text{max}}$), the number of TBSs ($n_{T}$), and the number of AVs ($M$). These combinations are detailed in Table III.
Instance $\boldsymbol{v_{\text{min}}}$ $\boldsymbol{v_{\text{max}}}$ $\boldsymbol{n_{T}}$ $\boldsymbol{M}$
I-(20,30,20,20) 20 m/s 30 m/s 20 20
I-(25,35,20,20) 25 m/s 35 m/s 20 20
I-(20,30,10,20) 20 m/s 30 m/s 10 20
I-(20,30,20,50) 20 m/s 30 m/s 20 50
I-(30,40,20,20) 30 m/s 40 m/s 20 20
Table III: Test Instances with Parameters.
@@ V-B Baselines and Evaluation Metrics
To comprehensively evaluate the performance of MO-DDQN-envelope, we implement the followsing baseline algorithms for comparison: MO-DQN, MO-DDQN (as discussed in Section IV-A), MO-dueling-DDQN, and MO-PPO. MO-dueling-DDQN and MO-PPO are variations of MO-DDQN, utilizing single-policy multi-objective dueling DDQN and Proximal Policy Optimization (PPO) algorithms, respectively, instead of DDQN. These algorithms, specifically developed for performance evaluation, are extensions of single-policy RL methods originally proposed in [42] and [43], respectively. Additionally, we incorporate several state-of-the-art algorithms from MORL-Baselines [40] for comparison. These include Scalarized Q-learning for single-policy MORL (MO-Q) [44], Scalarized DQN for single-policy DQN (MO-DQN) [1], Action Branching Architectures with Dueling DDQN for MORL (MO-Dueling-DDQN) [11], Multi-objective exploration for proximal policy optimization (MO-PPO) [45], and Expected Utility Policy Gradient (EUPG) [46]. To evaluate the performance, we consider the following metrics. Assume episode $e$ ends on time step $T_{e}$. we define (1) total transportation reward: $R^{tran}_{e}=\mathbb{E}_{j\in M_{1}}[\sum_{t=1}^{T_{e}}r^{\mathrm{j,tran}}_{t}]$. (2) total telecommunication reward: $R^{tele}_{e}=\mathbb{E}_{j\in M_{1}}[\sum_{t=1}^{T_{e}}r^{\mathrm{j,tele}}_{t}]$. (3) collision rate: $\delta_{e}=1-\frac{T_{e}}{T_{hl}}$ (4) HOs Probability: $\xi_{e}=\mathbb{E}_{j\in M_{1}}[\xi^{j}_{T_{e}}]$. where $r^{\mathrm{j,tran}}_{t}$ and $r^{\mathrm{j,tele}}_{t}$ are instantaneous transportation and telecommunication rewards specified in equations (17) and (18), respectively. $T_{hl}$ represents the horizon limit of each episode. $\xi^{j}_{T_{e}}$ denotes HOs probability by the end of each episode ($T_{e}$’s step) defined in equation (18). Table IV evaluate total transportation reward (17), total telecommunication reward (18), total rewards defined by $R^{total}_{e}=R^{tele}_{e}+R^{tran}_{e}$ over MO-DQN, MO-DDQN, and MO-DDQN-Envelope.
Desire Velocity Model Number of AVs $M$
(m/s) $M$=10 $M$=20 $M$=30 $M$=40
15 MO-DQN 23.42/10.21/33.63 22.71/9.18/31.89 21.09/7.45/28.54 18.35/6.12/24.47
MO-DDQN 24.28/12.31/36.59 24.89/11.41/36.30 21.85/8.63/30.48 19.41/7.11/26.52
MO-DDQN-Envelope 25.76/13.45/39.21 23.74/10.03/33.77 23.16/9.86/33.02 20.83/7.89/28.72
20 MO-DQN 20.85/8.56/29.41 19.92/6.74/26.66 19.23/6.35/25.58 17.48/4.39/21.87
MO-DDQN 21.62/10.19/31.81 21.04/8.57/29.61 20.11/7.93/28.04 18.72/5.97/24.69
MO-DDQN-Envelope 22.49/11.45/33.94 21.98/9.92/31.90 20.87/8.72/29.59 19.61/6.18/25.79
25 MO-DQN 18.21/7.48/25.69 17.45/5.91/23.36 16.34/5.12/21.46 14.87/4.37/19.24
MO-DDQN 18.92/8.19/27.11 17.88/6.72/24.60 16.97/5.94/22.91 15.32/4.92/20.24
MO-DDQN-Envelope 19.64/9.01/28.65 18.37/7.51/25.88 17.04/6.52/23.56 16.01/5.38/21.39
30 MO-DQN 16.09/6.32/22.41 15.76/5.03/20.79 14.92/4.76/19.68 13.81/4.21/18.02
MO-DDQN 16.72/6.89/23.61 16.21/5.32/21.53 15.43/5.14/20.57 14.89/4.91/19.80
MO-DDQN-Envelope 17.35/7.45/24.80 16.88/5.97/22.85 15.97/5.63/21.60 14.12/4.49/18.61
Table IV: Evaluation performance of MO-DQN, MO-DDQN, and MO-DDQN-Envelope (total transportation reward / total telecommunication reward / total reward).